OpenStudy (anonymous):

Since x^2+8x-15 cant be factored perfectly, what's the next step?

OpenStudy (alexwee123):

quadratic formula?

OpenStudy (tiffanymak1996):

u can use the quadratic formula to find x just let the equation be 0 before u use it.

OpenStudy (helder_edwin):

u MUST try the quadratic formula: if \(ax^2+bx+c=0\) then \[ \large x=\frac{-b\pm\sqrt{b^2-4ac}}{2a} \]

OpenStudy (tiffanymak1996):

a=1; b=8 c=-15

OpenStudy (helder_edwin):

yes

OpenStudy (tiffanymak1996):

and then you substitute a b c into the formula and you'll have x!

OpenStudy (helder_edwin):

yes. sorry, busy elsewhere.